OpenStudy (anonymous):

Simplify the number using the imaginary unit i Square root of -28

OpenStudy (anonymous):

\[\sqrt{-28}=\sqrt{-1*28}=\sqrt{28\iota^2}=\sqrt{28}\iota \]

OpenStudy (anonymous):

Thank u